OpenStudy (anonymous):

A pill is in the shape of a right circular cylinder with hemisphere on each end The cylinder and the hemispheres have the same radius. Find the volume of a pill with a radius of 7 millimeters and with its cylinder portion having a length of 16 millimeter.

OpenStudy (ybarrap):

The total volume would be the sum of the volumes of the cylinder and two hemispheres (i.e. one sphere): $$ V=\pi r^2h+\frac{4}{3}\pi r^3$$ Where r is radius of sphere and h is the length of the cylinder.